OpenStudy (anonymous):

For f(x)=2/3(x-5) determine f(-4)

OpenStudy (anonymous):

Plug it iiiin, guurl.

OpenStudy (anonymous):

\[\huge f(x) = \frac{2}{3} (x-5)\]? substitute x = -4

OpenStudy (anonymous):

aka it would be f(x)=2/3(-4-5)